OpenStudy (anonymous):

What is the approximate volume of a tree trunk if the trunk is 13 feet tall with a circumference of 4.5 feet? A. 33.09 ft3 B. 21.17 ft3 C. 17.16 ft3 D. 27.14 ft3

OpenStudy (joannablackwelder):

What kind of shape do you think the tree trunk would have?

jhonyy9 (jhonyy9):

so you know that the circumference of a circle is 2pi*r ,so this mean that 4,5 = 2pi*r 4,5 = 2*3,14*r 4,5= 6,28*r r = 4,5/6,28 = 0,716 feet r = 0,716 feet so than you can calculi area of base ,what will be pi*r^2 area of base = 3,14*(0,716)^2 = 3,14*0,512 = 1,607 feet^2 A = 1,607 feet ^2 Volume equal area of base time tall of the tree trunk V = 1,607*13 = 20,891 feet^3 depend how was rounded these results - so from the above choices the B. is right sure